# Graph Transformers without Positional Encodings

### v2

@@ Appendix A Proof of Proposition 1
If $SPD^{g_{1}}[i,j]=SPD^{g_{2}}[p,q]$, we simply set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=SPD^{g_{1}}[i,j]$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=SPD^{g_{2}}[p,q]$.
If $SPD^{g}[i,j]\neq SPD^{g}[p,q]$ and ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}\neq{v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, we are done. If however, ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}={v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, choose an index $r$ such that $\lambda^{g_{1}}_{r}\neq\lambda^{g_{2}}_{r}$ (which exists necessarily since otherwise $SPD^{g_{1}}[i,j]=SPD^{g_{2}}[p,q]$ due to 12). We can then perturb $\theta_{r}$ such that ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}$ becomes unequal to ${v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, in which case we again can set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=SPD^{g_{1}}[i,j]$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=SPD^{g_{2}}[p,q]$. Similarly to Case 2 above, we are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 13 for other graphs and node-pairs in $\mathbb{G}_{n}$.
@@
If $SPD^{g_{1}}[i,j]=SPD^{g_{2}}[p,q]$, we simply set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=SPD^{g_{1}}[i,j]$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=SPD^{g_{2}}[p,q]$.
If $SPD^{g}[i,j]\neq SPD^{g}[p,q]$ and ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}\neq{v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, we are done. If however, ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}={v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, choose an index $r$ such that $v^{g_{1}}_{ij}[r]\neq v^{g_{2}}_{pq}[r]$. We can then perturb $\theta_{r}$ such that ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}$ becomes unequal to ${v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, in which case we again can set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=SPD^{g_{1}}[i,j]$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=SPD^{g_{2}}[p,q]$. Similarly to Case 2 and 3 above, we are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 13 for other graphs and node-pairs in $\mathbb{G}_{n}$ since the set of such constraints is finite and bounded.
Thus, we have proven that the continuous function $\phi_{2}(\lambda)=\sum_{k=0}^{m}\theta_{k}(1-\lambda)^{k}$ for some $m\in\mathbb{Z}$ can be chosen to ensure the continuity of $\phi_{1}$ defined in 13. Both functions can hence be approximated by MLPs to an arbitrary accuracy due to Hornik et al. [1989].
@@ B.1 Description of Datasets
Dataset #graphs #nodes #edges Directed Prediction Task
@@
MNIST 70,000 $\sim$70.6 $\sim$564.5 No 10-class Graph classification
PATTERN 14,000 $\sim$118.9 $\sim$3039.3 No Binary node classification
CLUSTER 12,000 $\sim$117.2 $\sim$2150.9 No 6-class Node classification
Peptides-func 15,535 $\sim$150.9 $\sim$307.3 No 10-label Graph classification
Peptides-struct 15,535 $\sim$150.9 $\sim$307.3 No 11-task Graph regression
@@ B.2 Hyperparameters
Due to a limited training budget we could not perform a large scale hyperparameter search and mostly borrowed configurations from Ma et al. [2023] and Rampášek et al. [2022], keeping in line with the parameter budgets, optimizer (AdamW with $betas=(0.9,0.99)$ and $eps=1e-8$) and learning rate schedule (warmup+cosine annealing) used therein. However, our experiments show that due to a very different attention mechanism, our method benefits from deeper and wider models, and lower dropout rates. More details are included in Table 4 and 5.
@@
Hyperparameter ZINC MNIST PATTERN CLUSTER
# Transformer Layers 12 3 10 16
# Heads 8 8 8 8
Node feature embedding 128 - - -
Edge feature embedding 128 - - -
Hidden dim 80 72 96 80
$\phi_{1}$ Hidden dim 80 72 104 80
$\phi_{2}$ Hidden dim 80 72 104 72
Dropout 0 0 0 0
Attention dropout 0.2 0.1 0.1 0.1
Graph pooling sum mean - -
Batch size 128 512 32 32
Learning rate 0.001 0.001 0.0005 0.0005
# Epochs 2000 200 100 100
# Warmup epochs 50 5 5 5
Weight decay 1e-5 1e-5 1e-5 1e-5
# Parameters 496713 97360 504885 507590
@@
Hidden dim 96 96
$\phi_{1}$ Hidden dim 80 80
$\phi_{2}$ Hidden dim 80 80
@@
# Parameters 509406 509503
B.3 Hardware resources
All experiments were carried out on an AWS ml.g5.2xlarge Sagemaker training instance with 32 GB of system RAM and a single NVIDIA A10G GPU with 24GB of VRAM.

### v3

@@ A.1 Proposition 1
If $f(SPD^{g_{1}}[i,j])=f(SPD^{g_{2}}[p,q])$, we simply set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=f(SPD^{g_{1}}[i,j])$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=f(SPD^{g_{2}}[p,q])$.
If $f(SPD^{g}[i,j)]\neq f(SPD^{g}[p,q])$ and ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}\neq{v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, we are done. If however, ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}={v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, choose an index $r$ such that $\lambda^{g_{1}}_{r}\neq\lambda^{g_{2}}_{r}$ (which exists necessarily since otherwise $SPD^{g_{1}}[i,j]=SPD^{g_{2}}[p,q]$ due to 15). We can then perturb $\theta_{r}$ such that ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}$ becomes unequal to ${v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, in which case we again can set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=f(SPD^{g_{1}}[i,j])$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=f(SPD^{g_{2}}[p,q])$. Similarly to Case 2 above, we are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 16 for other graphs and node-pairs in $\mathbb{G}_{n}$.
@@
If $f(SPD^{g_{1}}[i,j])=f(SPD^{g_{2}}[p,q])$, we simply set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=f(SPD^{g_{1}}[i,j])$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=f(SPD^{g_{2}}[p,q])$.
If $f(SPD^{g}[i,j])\neq f(SPD^{g}[p,q])$ and ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}\neq{v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, we are done. If however, ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}={v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, choose an index $r$ such that $v^{g_{1}}_{ij}[r]\neq v^{g_{2}}_{pq}[r]$. We can then perturb $\theta_{r}$ such that ${v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2}$ becomes unequal to ${v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2}$, in which case we again can set $\phi_{1}({v^{g_{1}}_{ij}}^{T}\phi^{g_{1}}_{2})=f(SPD^{g_{1}}[i,j])$ and $\phi_{1}({v^{g_{2}}_{pq}}^{T}\phi^{g_{2}}_{2})=f(SPD^{g_{2}}[p,q])$. Similarly to Case 2 and 3 above, we are assured that such a perturbation exists without violating the continuity constraints of $\phi_{1}$ in 16 for other graphs and node-pairs in $\mathbb{G}_{n}$.
Thus, we have proven that the continuous function $\phi_{2}(\lambda)=\sum_{k=0}^{m}\theta_{k}(1-\lambda)^{k}$ for some $m\in\mathbb{Z}$ can be chosen to ensure the continuity of $\phi_{1}$ defined in 16. Both functions can hence be approximated by MLPs to an arbitrary accuracy due to Hornik et al. [1989].
@@ B.1 Description of Datasets
Dataset #graphs #nodes #edges Directed Level Task
@@
MNIST 70,000 $\sim$70.6 $\sim$564.5 Yes Graph 10-Class Classification
CIFAR10 60,000 $\sim$117.6 $\sim$941.2 Yes Graph 10-Class Classification
PATTERN 14,000 $\sim$118.9 $\sim$3039.3 No Node Binary Classification
CLUSTER 12,000 $\sim$117.2 $\sim$2150.9 No Node 6-Class Classification
Peptides-func 15,535 $\sim$150.9 $\sim$307.3 No Graph 10-Class Multilabel
Classification
Peptides-struct 15,535 $\sim$150.9 $\sim$307.3 No Graph 11-Task Regression
@@ B.2 Hyperparameters
Due to a limited training budget we could not perform a large scale hyperparameter search and mostly borrowed configurations from Ma et al. [2023] and Rampášek et al. [2022], keeping in line with the parameter budgets, optimizer (AdamW with $betas=(0.9,0.99)$ and $eps=1e-8$) and learning rate schedule (warmup+cosine annealing) used therein. Notably, due to the simple attention mechanism of Eigenformer (without feature attention), we could use deeper/wider models for training. Further details are included in Table 5 and 6 (SSR = Signed-square-root. For the notation A/B, A=metric without feature attention and B=metric with feature attention).
@@
Hyperparameter ZINC MNIST CIFAR10 PATTERN CLUSTER
# Transformer Layers 12 3 3 10 16
# Heads 8 8 8 8 8
$\psi$ SSR SSR Identity SSR SSR
Node feature embedding 128 - - - -
Edge feature embedding 128 - - - -
Hidden dim. 72/56 64/56 64/56 64/64 56/56
$\phi_{1}$ Hidden dim. 28/28 28/24 28/24 36/28 28/24
$\phi_{2}$ Hidden dim. 28/28 28/24 28/24 36/28 28/24
Dropout 0 0 0 0 0
Attention dropout 0.2 0.2 0.2 0.5 0.5
Graph pooling sum mean mean - -
Batch size 128/512 512 512 16/24 16
Learning rate 0.001 0.001 0.001 0.0005 0.0005
# Epochs 2000 200 200 100 100
# Warmup epochs 50 5 5 5 5
Weight decay 1e-5 1e-5 1e-5 1e-5 1e-5
# Parameters (w/o feat) 509849 106746 106874 476929 486006
# Parameters (w/ feat) 479481 110234 110346 472321 479734
# Epoch time (s) 21/23 198/222 793/840 571/619 550/582
# Accelerator (NVIDIA) A10G A10G A10G A10G A10G
# Peak memory (GB) 19.44/20.54 19.95/19.19 20.39/19.06 16.38/21.17 17.9/17.53
@@
$\psi$ SSR SSR
Hidden dim 80/64 80/64
$\phi_{1}$ Hidden dim 32/28 32/28
$\phi_{2}$ Hidden dim 32/28 32/28
@@
# Parameters (w/o feat) 510570 510651
# Parameters (w/ feat) 510651 478779
# Epoch time (s) 1740/2736 2541/2700
# Accelerator (NVIDIA) H100 H100
# Peak memory (GB) 60.76/75.76 70.77/66.69